byeeee:

Sandy is observing the velocity of a runner at different times. After one hour, the velocity of the runner is 4 km/h. After two hours, the velocity of the runner is 2 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the runner at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equations obtained in Part A for the first 4 hours? (5 points)

byeeee:

@Angle

Angle:

Ok so standard form is ax + by = c right?

byeeee:

yes..

Angle:

do you have any ideas on how to make the given information into any type of equation? it doesn't have to be the type they suggested for now

Angle:

for example, the slope intercept form: y = mx + b or the point-slope form: y = a(x - h) + k ?

byeeee:

nope not really the velocity and the fact its km its just confusing to me

Angle:

mm yeah, it's a lot of words and a lot of extra information I don't think we need to pay that much attention to the units, because 4 km/h and 2 km/h is using the same[i/] units

Angle:

whoops, formatted wrong

Angle:

in any case the information we need to use to get equations is: at hour 1, they went at a speed of 4 at hour 2, they went at a speed of 2

byeeee:

yeah alright

Angle:



byeeee:

ok...

Angle:

so the idea is that y = m(x - h) + k m = slope (h, k) = (x, y) point on the line

Angle:

slope is: "how much does it go up or down, if you move to the right 1 unit"? in this case, we are going from 4 to 2 if we move to the right by one spot

byeeee:

yes

Angle:

so what do you think the slope is?

byeeee:

(1,-3)

Angle:

hmmm let's try this this way slope = \(\large \frac{d - e}{f - g}\) \(d = 2\) \(e = 4\) \(f = 2\) \(g = 1\)

byeeee:

2-4 over 2-1

Angle:

mhm and that can be simplified to?

byeeee:

-1 idk that kinda confused me

Angle:

\(\large \frac{2-4}{2-1} = \frac{-2}{1} = -2\)

byeeee:

oh

Angle:



byeeee:

ohh i was going down 3 im silly

Angle:

"y = m(x - h) + k m = slope (h, k) = (x, y) point on the line" so m = slope = -2 then a point on the line is (1, 4) plug this all in and: y = (-2)*(x - 1) + 4

byeeee:

oh

byeeee:

so wait then for the way they are asking would a good equation be y = –2x + 6

Angle:

close! they want it in "standard form" so the x and y variables on the same side

Angle:

>just add 2x to both sides to move it

byeeee:

hm

Angle:

so we want it to look like this 2x + y = 6

byeeee:

right

Angle:

so that's part 1

Angle:

*part A

byeeee:

yep now to show how to graph is

byeeee:

it

Angle:

I graphed part of it, you just need to extend it to include hours 3 and 4

byeeee:

i think i got it

Angle:

great :)

byeeee:

thank you so much